# Probabilistic-based Feature Embedding of 4-D Light Fields for Compressive Imaging and Denoising

### v1

@@ 6.2.1 Comparisons on Simulated CMs
We compared our method, namely Ours, with three state-of-the-art deep learning-based LF reconstruction methods from CMs, i.e., Inagaki et al. (2018), Vadathya et al. (2019), and Guo et al. (2021). For a fair comparison, all methods were re-trained on the same training datasets with officially released codes and recommended configurations. We conducted three tasks on the Lytro dataset, i.e., reconstructing $7\times 7$ LFs with the input measurement number $S=1$, $S=2$, and $S=4$, respectively. We also conducted three tasks on the HCI+Inria dataset, i.e., reconstructing $5\times 5$ LFs with the input measurement number $S=1$, $2$, and $4$, respectively.
@@
  • Ours achieves better performance with a smaller model size than all compared methods on most tasks, which gives credit to our physical interpretable cycle-consistent framework and the probabilistic-based feature embedding strategy;
@@
  • Inagaki et al. (2018) has significantly lower PSNR and SSIM values than Ours on both Lytro and HCI+Inria datasets. The reason may be that Inagaki et al. (2018) simply employs 2-D convolutional filters that are unable to model the 4-D LF well;
@@
  • Vadathya et al. (2019) outperforms Ours under task $S=1$ on the Lytro dataset, which may be credited to its explicit use of geometry information. However, our method achieves much higher PSNR, i.e., about $5$ dB, than Vadathya et al. (2019) under tasks $S=2$ and $S=4$ on the Lytro dataset, because our probabilistic-based feature embedding strategy has the stronger representative ability to model the dimensional correlations among the LF with more than one CMs. Moreover, our method significantly outperforms Vadathya et al. (2019) under all tasks on the HCI+Inria dataset. We believe the reason is that LFs contained in this dataset have relatively large disparity, resulting in heavily blurred CMs, and thus, Vadathya et al. (2019) could have difficulties predicting a high-quality central SAI for view reconstruction; and
@@
  • Ours outperforms Guo et al. (2021) on both Lytro and HCI+Inria datasets. The possible reason is that Guo et al. (2021) employs the SAS convolutional layer, which combines spatial and angular dimensions in an empirical manner, while our method can adaptively learn the fusion of the spatial-angular features.
@@
Qualitative comparisons. Fig. 6 and Fig. 7 show visual comparisons of reconstructed LFs from different methods. We can observe that Ours produces better visual results than all the compared methods. Specifically, Inagaki et al. (2018) and Vadathya et al. (2019) show blurry effects, and lose high-frequency details at texture regions, while Ours can reconstruct sharp details at both texture and smooth regions. Besides, Guo et al. (2021) shows distortions at occlusion boundaries, while our method can reconstruct clearer and more accurate structures.
@@
Figure 9: Comparisons of our method against Inagaki et al. (2018) and Guo et al. (2021) on real CMs. The PSNR/SSIM values are calculated between the reconstructed LF and the pinhole images.
@@
Comparisons of the LF parallax structure. The quality of the parallax structure is one of the most important criteria for evaluating the reconstructed LF. To compare the ability to preserve the LF parallax structure, we visualized EPIs of the reconstructed LFs using different methods. As shown in Fig. 6 and Fig. 7, we observe that Ours shows clearer and sharper linear structures than the compared methods, which demonstrates better preservation of the parallax structures of reconstructed LFs. Moreover, the accuracy of the depth map estimated from the reconstructed LF reflects how well the parallax structure is preserved to some extent. Therefore, we compared the depth maps estimated from different methods using the same depth estimation method Chen et al. (2018b) and the ground truth depth maps. As shown in Fig. 8, we can see that the depth maps of Ours are more similar to those of the ground truth at both smooth regions and occlusion boundaries. This means that our method can better preserve the parallax structure than other methods
@@ 6.2.2 Comparisons on Real CMs
To demonstrate the ability of our method on real CMs, we adopted the real data captured by a coded aperture camera built by Inagaki et al. Inagaki et al. (2018). The real data contains $2$ CMs, $2$ masks that are used for capturing the CMs, and the captured $5\times 5$ pinhole images as the ground truth. We fixed the weights of $\mathcal{P}$ to the mask values, and re-trained the CR-Net under the task $S=2$ on the HCI+Inria dataset. Then, we tested the trained model with the real CMs, and compared the reconstructed LF by our method against that of Inagaki et al. (2018) and Guo et al. (2021). The results are shown in Fig. 9, where it can be observed that the reconstructed LF from our method is closer to the ground-truth one, while the results from Inagaki et al. (2018) show blurring artifacts at the occlusion boundaries, demonstrating the advantage of our method on real CMs.
@@ 6.2.3 Ablation Study
The number of stages and feature embedding units for the CR-Net. To investigate the influence of the number of stages and feature embedding units on the performance, we compared the results of CR-Net using $T=2$, $4$ and $6$ and $J=6$, $8$, and $10$. As demonstrated in Figure 10 and Table 3, the PSNR/SSIM values gradually improve from $2$ to $6$ stages for Ours_CR-Net and Ours_Template. Moreover, $8$ feature embedding units yield better reconstruction quality. Consequently, for superior LF reconstruction quality, we selected $6$ stages and $8$ feature embedding units in our CR-Net.
@@
Effectiveness of modeling probabilistic space. In Fig. 12, we visualized the finally learned architecture for the probabilistic-based feature embedding module, where it can be seen that both the layer-wise feature extraction and network-level feature aggregation are adaptively constructed based on the probabilistic modeling. Furthermore, to evaluate the effectiveness of modeling probabilistic space, we quantitatively compared the performance of the final learned CR-Net against the template network. As shown in Fig. 10, it can be observed that the final CR-Net consistently achieves higher performance with a smaller model size compared with the template network under different numbers of stages, which demonstrates the effectiveness of modeling probabilistic space for space-angular fusion in our method.
Effectiveness of the probabilistic-based feature embedding module. To verify the effectiveness of the probabilistic-based feature embedding module $\mathcal{D}(\cdot)$, we replaced $\mathcal{D}(\cdot)$ in each stage with either a stack of 3-D convolution layers or a stack of SAS layers while keeping a comparable model size to our CR-Net. The resulted networks are denoted as Ours_Conv3D and Ours_SAS, respectively, and their performance was compared with Ours_CR-Net in Table 4. The results indicate that replacing the probabilistic-based feature embedding module with 3-D convolution or SAS layers leads to obvious performance degradation, which further demonstrates the advantage of the proposed probabilistic-based feature embedding strategy.
Effectiveness of sharing weights in projection layers. We also verified the effectiveness of sharing weights of the projection layers across different stages. As shown in Table 4, the performance of the model without sharing the weights of the projection layers, denoted as w/o sharing $\mathcal{P}(\cdot)$, is lower than Ours_CR-Net. The reason is that sharing weights guarantees that the projection layers across all stages correspond to the same physical imaging process. As the pseudo-CMs are produced under the same projection, the residual between the pseudo-CMs and the input CMs is consistent and can be minimized progressively.
@@
Table 4: Effectiveness of the probabilistic-based feature embedding module and sharing the weights of projection layers in the CR-Net. The results on task $S=2$ over the HCI+Inria dataset are provided for comparisons.
@@
Ours_Conv3D(1) 40.42 0.963 1.78
Ours_Conv3D(2) 41.34 0.966 2.67
Ours_SAS 41.11 0.966 1.78
Ours_CR-Net 42.49 0.973 1.75
Table 5: Quantitative comparisons (PSNR/SSIM) of different methods over the test LFs in Stanford Archive and HCI+Inria datasets under three denoising tasks. The number of network parameters (M) of different methods are also provided. “-” indicates that the method is not a deep learning-based method.
@@
Stanford Archive Ren et al. (2021) 37.22/0.971 33.69/0.941 29.38/0.876 2.23
Alain and Smolic (2017) 38.25/0.927 33.27/0.841 25.11/0.642 -
Tomita et al. (2022) 39.40/0.916 36.67/0.861 32.92/0.742 -
Chen et al. (2018a) 39.34/0.956 36.94/0.939 33.65/0.894 1.25
Guo et al. (2021) 42.77/0.973 39.98/0.958 35.96/0.920 2.22
Ours 43.20/0.978 40.59/0.964 36.72/0.931 2.39
HCI+Inria Ren et al. (2021) 38.05/0.976 34.71/0.955 30.65/0.919 2.23
Alain and Smolic (2017) 40.83/0.965 37.36/0.936 31.94/0.860 -
Tomita et al. (2022) 38.95/0.940 35.93/0.907 31.96/0.831 -
Chen et al. (2018a) 36.89/0.933 33.92/0.896 30.19/0.813 1.22
Guo et al. (2021) 42.10/0.971 39.53/0.955 35.14/0.899 2.22
Ours 43.21/0.976 40.50/0.962 36.24/0.918 2.39
@@ 6.3 Evaluation on 4-D LF Denoising
We compared our method against several state-of-the-art methods to evaluate our extension on LF denoising, including a local similarity-based LF denoising method denoted by LFBM5D Alain and Smolic (2017), a short-time DFT approach denoted by Mviden Tomita et al. (2022), a deep learning-based LF denoising method by APA Chen et al. (2018a), a deep unrolling-based LF denoising method by DSAR Guo et al. (2021), and a deep learning-based single image denoising method DeamNet Ren et al. (2021) which was applied on each SAI of the input LF as a baseline.
The same noise synthesis and preprocessing protocol as APA were used in our experiment, i.e., adding zero-mean Gaussian noise with the standard variance $\sigma$ varying in the range of 10, 20, and 50 to generate the noisy LF images. We use the same training and test datasets with a narrow baseline from Stanford Lytro Light Field Archive Raj et al. as APA, extracting central $8\times 8$ SAIs and converting to grayscale for each LF. In addition, we performed experimental validation on the wider baseline HCI+Inria dataset, in which central $7\times 7$ SAIs were used, $55$ and $6$ LF images were used for training and testing, respectively. We trained all the comparison methods for each noise level, and evaluated each model with the matched noise level on test data.
Table 6: Comparisons of the running time (in seconds) of different methods for HCI+Inria denoising tasks.
Algorithms Ren et al. (2021) Alain and Smolic (2017) Tomita et al. (2022) Chen et al. (2018a) Guo et al. (2021) Ours
HCI+Inria 2.35 79.65 118.51 17.11 14.72 11.63
@@ 6.3.1 Quantitative Comparison
The average PSNR and SSIM between the denoised LFs and ground-truth ones were used to evaluate different methods quantitatively. From Table 5, it can be observed that Ours outperforms the second-best method, i.e., DSAR, up to $0.5dB$ on Stanford Archive for all three noise levels, and up to $1.1dB$ on HCI+Inria for all three noise levels, validating the advantage of our network. Besides, the performance advantage of Ours is more obvious on wider baseline LFs, validating our method’s ability to process large baseline dataset.
6.3.2 Visual Comparison
Fig. 13 and Fig. 14 present visual comparisons of denoised LFs obtained through different methods across three noise levels $\sigma=10,20$ and $50$. The results indicate that DeamNet, LFBM5D, and Mviden fail to preserve high-frequency details, such as the texture regions of flowers, and generate severe distortions, such as the window frames. Additionally, DeamNet produces hole artifacts in the denoised LFs. While APA yields comparatively better visual results than LFBM5D and DeamNet, it still exhibits blurring artifacts at occlusion boundaries. DASR achieves relatively good visual results, but the details in textured regions and occlusion boundaries are not as sharp as those produced by our method. Visual comparison of the EPIs extracted from the denoised LFs indicates that our method can preserve sharper linear structures, thereby validating its advantage in preserving LF parallax structures during denoising.
Furthermore, we compared the estimated depth maps obtained through different methods using the same depth estimation method Chen et al. (2018b). As shown in Figure 15, the depth maps generated by our method are more similar to the ground truth at both smooth regions and occlusion boundaries, indicating that our method can better preserve the parallax structure than the other methods.

### v3

@@ 6.2.1 Comparisons on Simulated CMs
We compared our CR-Net with three state-of-the-art deep learning-based LF reconstruction methods from CMs, i.e., Inagaki et al. (2018), Vadathya et al. (2020), and Guo et al. (2022). For a fair comparison, all methods were re-trained on the same training datasets with officially released codes and recommended configurations. We conducted three tasks on the Lytro dataset, i.e., reconstructing $7\times 7$ LFs with the input measurement number $S=1$, $S=2$, and $S=4$, respectively. We also conducted three tasks on the HCI+Inria dataset, i.e., reconstructing $5\times 5$ LFs with the input measurement number $S=1$, $2$, and $4$, respectively.
@@
  • CR-Net achieves better performance with a smaller model size than all compared methods on most tasks, which gives credit to our physical interpretable cycle-consistent framework and the probabilistic-based feature embedding strategy;
@@
  • Inagaki et al. (2018) has significantly lower PSNR and SSIM values than CR-Net on both Lytro and HCI+Inria datasets. The reason may be that Inagaki et al. (2018) simply employs 2-D convolutional filters that are unable to model the 4-D LF well;
@@
  • Vadathya et al. (2020) outperforms CR-Net under task $S=1$ on the Lytro dataset, which may be credited to its explicit use of geometry information. However, our CR-Net achieves much higher PSNR, i.e., about $5$ dB, than Vadathya et al. (2020) under tasks $S=2$ and $S=4$ on the Lytro dataset, because our probabilistic-based feature embedding strategy has the stronger representative ability to model the dimensional correlations among the LF with more than one CMs. Moreover, the proposed CR-Net significantly outperforms Vadathya et al. (2020) under all tasks on the HCI+Inria dataset. We believe the reason is that LFs contained in this dataset have relatively large disparity, resulting in heavily blurred CMs, and thus, Vadathya et al. (2020) could have difficulties predicting a high-quality central SAI for view reconstruction; and
@@
  • CR-Net outperforms Guo et al. (2022) on both Lytro and HCI+Inria datasets. The possible reason is that Guo et al. (2022) employs the SAS convolutional layer, which combines spatial and angular dimensions in an empirical manner, while our method can adaptively learn the fusion of the spatial-angular features.
@@
Qualitative comparisons. Fig. 6 and Fig. 7 show visual comparisons of reconstructed LFs from different methods. We can observe that CR-Net produces better visual results than all the compared methods. Specifically, Inagaki et al. (2018) and Vadathya et al. (2020) show blurry effects, and lose high-frequency details at texture regions, while our CR-Net can reconstruct sharp details at both texture and smooth regions. Besides, Guo et al. (2022) shows distortions at occlusion boundaries, while our CR-Net can reconstruct clearer and more accurate structures.
@@
Figure 9: Comparisons of our CR-Net against Inagaki et al. (2018) and Guo et al. (2022) on real CMs. The PSNR/SSIM values are calculated between the reconstructed LF and the pinhole images.
@@
Comparisons of the LF parallax structure. The quality of the parallax structure is one of the most important criteria for evaluating the reconstructed LF. To compare the ability to preserve the LF parallax structure, we visualized EPIs of the reconstructed LFs using different methods. As shown in Fig. 6 and Fig. 7, we observe that the proposed CR-Net shows clearer and sharper linear structures than the compared methods, which demonstrates better preservation of the parallax structures of reconstructed LFs. Moreover, the accuracy of the depth map estimated from the reconstructed LF reflects how well the parallax structure is preserved to some extent. Therefore, we compared the depth maps estimated from different methods using the same depth estimation method Chen et al. (2018b) and the ground truth depth maps. As shown in Fig. 8, we can see that the depth maps of CR-Net are more similar to those of the ground truth at both smooth regions and occlusion boundaries. This means that our method can better preserve the parallax structure than other methods.
@@ 6.2.2 Comparisons on Real CMs
To demonstrate the ability of our method on real CMs, we adopted the real data captured by a coded aperture camera built by Inagaki et al. Inagaki et al. (2018). The real data contains $2$ CMs, $2$ masks that are used for capturing the CMs, and the captured $5\times 5$ pinhole images as the ground truth. We fixed the weights of $\mathcal{P}$ to the mask values, and re-trained the CR-Net under the task $S=2$ on the HCI+Inria dataset. Then, we tested the trained model with the real CMs, and compared the reconstructed LF by our CR-Net against that of Inagaki et al. (2018) and Guo et al. (2022). The results are shown in Fig. 9, where it can be observed that the reconstructed LF from our CR-Net is closer to the ground-truth one, while the results from Inagaki et al. (2018) show blurring artifacts at the occlusion boundaries, demonstrating the advantage of our CR-Net on real CMs.
@@ 6.2.3 Ablation Study
The number of stages and feature embedding units for the CR-Net. To investigate the influence of the number of stages and feature embedding units on the performance, we compared the results of CR-Net using $T=2$, $4$ and $6$ and $J=6$, $8$, and $10$. As demonstrated in Fig. 10 and Table 3, the PSNR/SSIM values gradually improve from $2$ to $6$ stages for our CR-Net and the template network. Moreover, $8$ feature embedding units yield better reconstruction quality. Consequently, for superior LF reconstruction quality, we selected $6$ stages and $8$ feature embedding units in our CR-Net.
@@
Effectiveness of modeling probabilistic space. In Fig. 12, we visualized the finally learned architecture for the probabilistic-based feature embedding module, where it can be seen that both the layer-wise feature extraction and network-level feature aggregation are adaptively constructed based on the probabilistic modeling. Furthermore, to evaluate the effectiveness of modeling probabilistic space, we quantitatively compared the performance of the final learned CR-Net against the template network. As shown in Fig. 10 and Table 4, it can be observed that the final CR-Net consistently achieves higher performance with a smaller model size compared with the template network under different numbers of stages, which demonstrates the effectiveness of modeling probabilistic space for space-angular fusion in our method.
Effectiveness of the probabilistic-based feature embedding module. To verify the effectiveness of the probabilistic-based feature embedding module $\mathcal{D}(\cdot)$, we replaced $\mathcal{D}(\cdot)$ in each stage with either a stack of 3-D convolution layers or a stack of SAS layers while keeping a comparable model size to our CR-Net. The resulting networks are denoted as CR-Conv3D and CR-SAS, respectively, and their performance was compared with our CR-Net in Table 4. The results indicate that replacing the probabilistic-based feature embedding module with 3-D convolution or SAS layers leads to obvious performance degradation, which further demonstrates the advantage of the proposed probabilistic-based feature embedding strategy.
Effectiveness of sharing weights in projection layers. We also verified the effectiveness of sharing weights of the projection layers across different stages. As shown in Table 4, the performance of the model without sharing the weights of the projection layers, denoted as w/o sharing $\mathcal{P}(\cdot)$, is lower than the CR-Net. The reason is that sharing weights guarantees that the projection layers across all stages correspond to the same physical imaging process. As the pseudo-CMs are produced under the same projection, the residual between the pseudo-CMs and the input CMs is consistent and can be minimized progressively.
@@
Table 4: Effectiveness of the probabilistic-based feature embedding module and sharing the weights of projection layers in the proposed CR-Net. The results on task $S=2$ over the HCI+Inria dataset are provided for comparisons.
@@
CR-Conv3D(1) 40.42 0.963 1.78
CR-Conv3D(2) 41.34 0.966 2.67
CR-SAS 41.11 0.966 1.78
CR-Template 42.12 0.972 1.83
CR-Net 42.49 0.973 1.75
Table 5: Quantitative comparisons (PSNR/SSIM) of different methods over the test LFs in Stanford Archive and HCI+Inria datasets under the denoising task of three noise levels. The number of network parameters (M) of different methods is also provided. “-” indicates that the method is not a deep learning-based method.
@@
Stanford Archive DeamNet Ren et al. (2021) 37.22/0.971 33.69/0.941 29.38/0.876 2.23
LFBM5D Alain and Smolic (2017) 38.25/0.927 33.27/0.841 25.11/0.642 -
Mviden Tomita et al. (2022) 39.40/0.916 36.67/0.861 32.92/0.742 -
APA Chen et al. (2018a) 39.34/0.956 36.94/0.939 33.65/0.894 1.25
DSAR Guo et al. (2022) 42.77/0.973 39.98/0.958 35.96/0.920 2.22
MSP Wang et al. (2023c) 42.37/0.974 39.67/0.962 36.06/0.929 1.25
DN-Net 43.20/0.978 40.59/0.964 36.72/0.931 2.39
HCI+Inria DeamNet Ren et al. (2021) 38.05/0.976 34.71/0.955 30.65/0.919 2.23
LFBM5D Alain and Smolic (2017) 40.83/0.965 37.36/0.936 31.94/0.860 -
Mviden Tomita et al. (2022) 38.95/0.940 35.93/0.907 31.96/0.831 -
APA Chen et al. (2018a) 36.89/0.933 33.92/0.896 30.19/0.813 1.22
DSAR Guo et al. (2022) 42.10/0.971 39.53/0.955 35.14/0.899 2.22
MSP Wang et al. (2023c) 41.51/0.972 38.80/0.956 35.31/0.913 1.23
DN-Net 43.21/0.976 40.50/0.962 36.24/0.918 2.39
@@ 6.3 Evaluation on 4-D LF Denoising
We compared our DN-Net against several state-of-the-art methods, including a local similarity-based LF denoising method denoted by LFBM5D Alain and Smolic (2017), a short-time DFT approach denoted by Mviden Tomita et al. (2022), two deep learning-based LF denoising methods named APA Chen et al. (2018a) and MSP Wang et al. (2023c), a deep unrolling-based LF denoising method DSAR Guo et al. (2022), and a deep learning-based single image denoising method DeamNet Ren et al. (2021) which was applied on each SAI of the input LF as a baseline.
The same noise synthesis and preprocessing protocol as APA were used in our experiment, i.e., adding zero-mean Gaussian noise with the standard variance $\sigma$ varying in the range of 10, 20, and 50 to generate the noisy LF images. We use the same training and test datasets with a narrow baseline from Stanford Lytro Light Field Archive Raj et al. as APA, extracting central $8\times 8$ SAIs and converting them to grayscale for each LF. In addition, we performed experimental validation on the wider baseline HCI+Inria dataset, in which central $7\times 7$ SAIs were used, $55$ and $6$ LF images were used for training and testing, respectively. We trained all the comparison methods for each noise level, and evaluated each model with the matched noise level on test data.
Table 6: Comparisons of the running time (in seconds) of different methods for denoising tasks on HCI+Inria.
Method Running Time
DeamNet Ren et al. (2021) 2.35
LFBM5D Alain and Smolic (2017) 79.65
Mviden Tomita et al. (2022) 118.51
APA Chen et al. (2018a) 17.11
DSAR Guo et al. (2022) 14.72
MSP Wang et al. (2023c) 4.16
Proposed DN-Net 11.63
@@ 6.3.1 Quantitative Comparison
The average PSNR and SSIM between the denoised LFs and ground-truth ones were used to evaluate different methods quantitatively. From Table 5, it can be observed that the proposed DN-Net outperforms the second-best method, i.e., DSAR, up to $0.5dB$ on Stanford Archive for all three noise levels, and up to $0.9dB$ on HCI+Inria for all three noise levels, validating the advantage of our network. Besides, the performance advantage of DN-Net is more obvious on wider baseline LFs, validating our method’s ability to process large baseline datasets. Additionally, Fig. 13 depicts the average PSNR at each SAI position of the denoised LFs by different methods. The performance of our DN-Net across different SAIs exhibits greater consistency and stability when compared to other methods. This demonstrates the robustness of our approach in maintaining high PSNR values for varying SAIs.
Figure 13: Comparison of the average PSNR for each SAI of denoised LFs produced by various methods. The PSNR values depicted are calculated as the average across the test set derived from the HCI+Inria dataset.
@@
6.3.2 Visual Comparison
Fig. 14 and Fig. 15 present visual comparisons of denoised LFs obtained through different methods across three noise levels $\sigma=10,20$ and $50$. The results indicate that DeamNet, LFBM5D, and Mviden fail to preserve high-frequency details, such as the texture regions of flowers, and generate severe distortions, such as the window frames. Additionally, DeamNet produces hole artifacts in the denoised LFs. While APA yields comparatively better visual results than LFBM5D and DeamNet, it still exhibits blurring artifacts at occlusion boundaries. DSAR and MSP achieve relatively good visual results, but the details in textured regions and occlusion boundaries are not as sharp as those produced by our DN-Net. Visual comparison of the EPIs extracted from the denoised LFs indicates that our method can preserve sharper linear structures, thereby validating its advantage in preserving LF parallax structures during denoising.
Furthermore, we compared the estimated depth maps obtained through different methods using the same depth estimation method Chen et al. (2018b). As shown in Fig. 16, the depth maps generated by our DN-Net are more similar to the ground truth at both smooth regions and occlusion boundaries, indicating that the proposed DN-Net can better preserve the parallax structure than the other methods.
Figure 17: Visual comparisons of different methods on real LF denoising. Our DN-Net trained on Stanford Archive dataset with $\sigma=20$ were selected. The zoomed selected regions and EPI are shown for better visualization
6.3.3 Comparisons on Real LF Denoising
When the lighting of the shooting scene is not sufficient, the light field data will be seriously affected by noise. Therefore, we used the low-light data collected by the Lytro Illum camera and verified the effectiveness of our DN-Net trained on the Stanford Archive dataset and with $\sigma=20$. From the experimental results in Fig. 17, we can observe that our algorithm can effectively remove noise while maintaining clear scene details better.